OpenStudy (anonymous):

Part 1: Find the perimeter of a rectangular object which has a length of square root 180 feet and a width of square root 125 feet. Part 2: Explain, in complete sentences, how you arrived at the answer and give the final solution in simplified radical form. Part 3: What type of object in your home or school might this be?

OpenStudy (anonymous):

do you add or multiple to find the perimeter?

jhonyy9 (jhonyy9):

- so you get the perimeter if you add 2 time width and 2 time lenght - so like P=2w +2l =2*sqrt125 + 2*sqrt180 =2*sqrt(5*25) +2*sqrt(9*20) = =2*5sqrt5 +2*3sqrt20 =10sqrt5 +6sqrt(5*4) =10sqrt5 +6*2sqrt5 = =10sqrt5 +12sqrt5 =22sqrt5 - so P=22sqrt5 feet